5) The bacterium Thermus thermophilus was first isolated from hot springs in Japan, living in an environment with an average temperature of 68° C (154 F). This temperature would kill you, but these organisms are adapted to it. One of the adaptations to this lifestyle includes DNA that is 70% G-C base pairs, one of the more extreme departures from the average 50%. The chromosome is about 2.13 x 10 bp in length, quite normal for a bacterial species. The restriction enzyme, Hae recognizes the degenerate sequence, 5-XGCGCTY-3, in which X can be any purine and Y can be any pyrimidine and the v represents the site of strand cleavage If you were to digest T thermophiles DNA with Hae what would be the average size of the fragments, and how many fragments would you expect to get? (10 pts)
Answer #1

Ans. DNA consists of 4 types of bases (A, C, T, G). A base at each position has the same probability of (1/4) at all positions- that is, probability of finding a “A” in [AAGCTT] at position 1 (first nucleotide) is (1/4); the same is also true for position 2 ; or any other base at any position. It is because the identity of a base at one position does not affect the identity of another base at different position.

So, the probability of finding 1 (any one) base out of total 4 type of base = 1/4.

The probability of finding a restriction site in a DNA molecule = (¼)n,

where n is the number of nucleotides in the restriction site.

Now,

#1. The probability of finding HaeII (n =6) restriction site = (1/4)6 = 1/ 4096 i.e. there is 1 restriction site every 4096 base pairs for dsDNA molecule.

Therefore, average size of fragments = 4096 base pairs

#2. No. of restriction fragments = Chromosome size / size of restriction fragment

                                                = (2.13 x 106 bp) / 4096 bp

                                                = 520